# Offline-online approximation of multiscale eigenvalue problems with random defects

### v1

@@ 2.3 Modified LOD (MLOD) method
The MLOD method: Find $(\tilde{\lambda}_{{H,\mathfrak{k}}},\tilde{u}_{{H,\mathfrak{k}}})\in\mathbb{R}\times V_{H}$ such that
@@ 2.4 The offline-online strategy
The equation (6) tells us that for each realization of $A$, we have to re-compute the correction operator $\mathscr{C}_{\mathfrak{k}}$, and consequently the LOD basis functions. This may be computationally very demanding although the computation of $\mathcal{C}_{\mathfrak{k}}$ can be parallelized. As a remedy, we use the offline-online strategy from [MV22], which we recall here to be self-contained.
@@
$\widetilde{a}_{\mathfrak{k}}(u_{{H,}\mathfrak{k}},v)=\sum\limits_{T\in\mathcal{T}_{H}}\widetilde{a}_{{T}}(u_{\scriptscriptstyle{H,}\mathfrak{k}},v_{H})$
@@
If $A=A_{\varepsilon}$, the local contribution $\widetilde{a}_{T}(\cdot,\cdot)$ is exactly the same for every mesh element $T\in\mathcal{T}_{H}$. This motivates to pre-compute $\widetilde{a}_{T}$ for certain defect configurations, but only one single fixed $T$, in the offline phase and then to recombine these quantities in a suitable manner for each realization of $A$.
@@ 2.4.1 Offline phase
This is the offline "basis" of coefficients where $A_{i}$ is constructed from $A_{\varepsilon}$ by considering only a single defect. Figure 2 shows the first four offline coefficients by using such a single defect on a single patch for the random checkerboard. This is similarly done for the random erasure by removing a single element from the perfectly periodic configuration.
@@
Next, we pre-compute and store the LOD stiffness matrix contributions
@@
for all our offline coefficients. We stress that the stiffnes matrix contribution for a fixed element $T$ is a coarse-scale object and inexpensive to store. At the same time, we can assemble the standard FE mass matrix.
@@ 2.4.2 Online Phase
$\widehat{a}_{\mathfrak{k}}=\sum\limits_{T\in\mathcal{T}_{H}}\widehat{a}_{\mathfrak{k},T}.$ (11)
@@
The Offline-Online MLOD method (OLOD method): Find $(\hat{\lambda}_{H,\mathfrak{k}},\hat{u}_{H,\mathfrak{k}})\in\mathbb{R}\times V_{H}$ such that,
@@
$\hat{u}^{\mathrm{ms}}_{H,\mathfrak{k}}=\hat{u}_{H,\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}_{H,\mathfrak{k}}\;\;\;\;\;\text{with }\;\;\;\;\widehat{\mathscr{C}}_{\mathfrak{k}}:=\sum\limits_{T\in\mathcal{T}_{H}}\sum\limits_{i=0}^{\mathcal{N}}\mu_{i}\mathscr{C}_{\mathfrak{k},T}(A_{i}).$
@@
$E^{2}_{T}:=\max\limits_{v\in V_{H}:v|_{T}}\frac{\norm{(A^{1/2}-A^{-1/2}\overline{A})\chi_{T}\nabla u-\sum\limits_{i=0}^{N}\mu_{i}(A^{1/2}-A^{-1/2}A_{i})\nabla(\mathscr{C}_{\mathfrak{k},T}(A_{i})v)}^{2}_{L^{2}(U_{\mathfrak{k}}(T))}}{\norm{v}^{2}_{A,T}}.$
Then, for any $v_{1},v_{2}\in V_{H}$ it holds that
|$\absolutevalue{(\widehat{a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}})(v_{1},v_{2})}\lesssim\mathfrak{k}^{d/2}\Big{(}\max\limits_{T\in\mathcal{T}_{H}}E_{T}\Big{)}\norm{v_{1}}_{\mathcal{A}}\norm{v_{2}}_{\mathcal{A}}.$ (13)
@@
$\begin{split}\norm{({\mathscr{C}_{\mathfrak{k}}}-\widehat{\mathscr{C}}_{\mathfrak{k}}){v}}^{2}_{\mathcal{A}}\lesssim\;\mathfrak{k}^{d}\Big{(}\max_{T\in\mathcal{T}_{H}}E_{T}\Big{)}^{2}\norm{v}^{2}_{\mathcal{A}}.\end{split}$ (14)
@@
$\begin{split}\widehat{a}_{\mathfrak{k}}(v,v)\geq(c_{1}-c_{2}g(\mathfrak{k})-\eta_{\mathfrak{k}})\norm{v}^{2}_{\mathcal{A}}\end{split}$ (15)
@@ 2.5 An alternate offline-online strategy
In this section, we define an enhanced offline-online strategy for random checkerboard-type coefficients, particularly when $Q=[0,1]^{d}$. Note that, if $Q$ is a proper subset of $[0,1]^{d}$, the sum constraint $\sum\mu_{i}=1$ is necessary to guarantee the correct value of the coefficient $A$ at points $x$ where defects can never occur. One example is the case of random erasure and points $x$ corresponding to the value $A_{\mathrm{per}}$ independent of the realization. However, in the random checkerboard case, the sum constraint $\sum\mu_{i}=1$ is merely used to guarantee uniqueness of the $\mu_{i}$. The alternate method we propose takes advantage of this by relaxing the sum constraint to be a real number $s$ that is adapted to the underlying distribution to obtain more accurate results. In the explained random checkerboard example, we use the Bernoulli distribution. However, other probability distributions on the interval $[\alpha,\beta]$, such as the uniform distribution, are possible in general as well. To this end, we consider the coefficient $A$ of the following form
$A(x,\omega)=\sum\limits_{j\in I}\chi_{\varepsilon(j+Q)}(x)\tilde{\beta}_{j}(\omega)$
@@
$A^{i}_{\text{harm}}=\frac{\mathcal{N}}{(\mathcal{N}-1)(1/\alpha)+(1)(1/\beta)}.$ (17)
@@
for $A_{\mathrm{per}}=\alpha$ and $B_{\mathrm{per}}=\beta-\alpha$ in the random checkerboard set-up, in which the underlying linear system of equations assigns $\tilde{\beta}_{i}$ value for each $\varepsilon$ block on the patch via the characteristic function $\chi_{\varepsilon(i+Q)}(x)$. Hence, we deduce
$\begin{split}\mu_{i}=\frac{\tilde{\beta}_{i}-\alpha s}{\beta-\alpha},\;\;j\geq 1,\hskip 56.9055pt\mu_{0}=\frac{s(\beta-\alpha)-\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}{\beta-\alpha}.\end{split}$ (18)
@@
$\begin{split}A^{\mu}_{\text{harm}}|_{T}&=\frac{\absolutevalue{T}}{\varepsilon}\frac{1}{\beta-\alpha}\Bigg{(}\frac{s(\beta-\alpha)-\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}{\mathcal{N}(1/\alpha)}+\frac{\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}{(\mathcal{N}-1)(1/\alpha)+(1/\beta)}\Bigg{)}\\ &=\frac{\absolutevalue{T}}{\beta-\alpha}\Bigg{(}\frac{s(\beta-\alpha)-\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}{\mathcal{N}\overline{A}}+\frac{\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}{(\mathcal{N}-1)\overline{A}+\overline{A}_{s}}\Bigg{)}\end{split}$
where $\overline{A}=\varepsilon/\alpha$ and $\overline{A}_{s}={\varepsilon}/{\beta}$. With this, we obtain for the error that
@@
$\begin{split}A_{\text{harm}}|_{T}-A^{\mu}_{\text{harm}}|_{T}&=\frac{\absolutevalue{T}}{\varepsilon}\Bigg{[}\frac{1}{\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i})^{-1}}-\frac{1}{\beta-\alpha}\Bigg{(}\frac{s(\beta-\alpha)-\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}{\mathcal{N}}\alpha+\frac{\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}{\frac{(\mathcal{N}-1)\beta+\alpha}{\alpha\beta}}\Bigg{)}\Bigg{]}\\ &=H(\tilde{\beta})-s\alpha+\frac{\alpha}{\beta-\alpha}\Bigg{(}\frac{\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}{\mathcal{N}}\Bigg{)}\mathcal{N}-\frac{\mathcal{N}^{2}\alpha\beta}{(\beta-\alpha)}\Bigg{[}\Bigg{(}\frac{\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}{\mathcal{N}}\Bigg{)}\frac{1}{(\mathcal{N}-1)\beta+\alpha}\Bigg{]}\\ &=H(\tilde{\beta})-s\alpha+\frac{M(\tilde{\beta}-\alpha s)}{(\beta-\alpha)}\frac{\alpha\mathcal{N}(\alpha-\beta)}{(\mathcal{N}-1)\beta+\alpha}\\ &=H(\tilde{\beta})-s\alpha-\alpha M(\tilde{\beta}-\alpha s)\frac{1}{(1-1/\mathcal{N})\beta+\alpha/\mathcal{N}}\end{split}$
@@
with $H(\tilde{\beta})=\mathcal{N}\Big{(}\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i})^{-1}\Big{)}^{-1}$ and $M(\tilde{\beta}-\alpha s)=\Big{(}{\sum\limits_{i=1}^{\mathcal{N}}(\tilde{\beta}_{i}-\alpha s)}\Big{)}/{\mathcal{N}}$. Our goal is to make this error as small as possible because this will imply a small consistency error. We consider the limit $\mathcal{N}\to 0$ which corresponds to vanishing periodicity length $\varepsilon\to 0$, and try to find $s$ such that the error vanishes in this limit. In other words, we find $s$ in such a way that in the homogenization limit, there is no consistency error and our offline-online strategy retrieves the homogenized coefficient. Denote
$L=\lim\limits_{\mathcal{N}\to\infty}H({\tilde{\beta}})\text{ and }M=\lim\limits_{\mathcal{N}\to\infty}M({\tilde{\beta}}).$
@@
Note that $L$ and $M$ can be deduced from the probability law of $\tilde{\beta}$ as certain expectations by the law of large numbers. Therefore, as $\mathcal{N}\to\infty$, we obtain
@@
$A_{\text{harm}}|_{T}-A^{\mu}_{\text{harm}}|_{T}=L-s\alpha-\frac{\alpha}{\beta}(M-\alpha s)=0.$
@@
$s=\frac{\beta L-\alpha M}{\alpha(\beta-\alpha)}.$
@@
$M=\alpha(1-p)+\beta p\;,\;\;\;\text{ and }\;\;\;L=\frac{\alpha\beta}{\beta+p(\alpha-\beta)}.$
@@
$s=1+\frac{p^{2}(\beta-\alpha)}{\beta+p(\alpha-\beta)}.$
Since the alternate $s$-value deviates from $1$ more the larger $p$ is, we expect that the influence of our new strategy is greater for larger $p$. This will be studied in more detail numerically in Section 5.

### v2

@@ 2.3 Modified LOD (MLOD) method
The MLOD problem: Find $(\tilde{\lambda}_{{H,\mathfrak{k}}},\tilde{u}_{{H,\mathfrak{k}}})\in\mathbb{R}\times V_{H}$ such that
@@ 2.4 The offline-online strategy
The equation (5) tells us that for each realization of $A$, we have to re-compute the correction operator $\mathscr{C}_{\mathfrak{k}}$, and consequently the LOD basis functions. This may be computationally very demanding although the computation of $\mathscr{C}_{\mathfrak{k}}$ can be parallelized. As a remedy, we use the offline-online strategy from [MV22], which we recall here to be self-contained.
@@
$\widetilde{a}_{\mathfrak{k}}(u_{{H,}\mathfrak{k}},v)=\sum\limits_{T\in\mathcal{T}_{H}}\widetilde{a}_{\mathfrak{k},{T}}(u_{\scriptscriptstyle{H,}\mathfrak{k}},v)$
@@
From this point onward, we assume that the mesh size $H$ is an integer multiple of the periodicity length $\varepsilon$. If $A=A_{\varepsilon}$, the local contribution $\widetilde{a}_{\mathfrak{k},T}(\cdot,\cdot)$ is exactly the same for every mesh element $T\in\mathcal{T}_{H}$. This motivates to pre-compute $\widetilde{a}_{\mathfrak{k},T}$ for certain defect configurations, but only one single fixed $T$, in the offline phase and then to recombine these quantities in a suitable manner for each realization of $A$.
@@ 2.4.1 Offline phase
This is the offline “basis” of coefficients where $A_{i}$ is constructed from $A_{\varepsilon}$ by considering only a single defect. Figure 2 shows the first four offline coefficients by using such a single defect on a single patch for the random checkerboard. This is similarly done for the random erasure by removing a single element from the perfectly periodic configuration.
@@
Next, for each offline coefficient, we pre-compute and store the LOD stiffness matrix contributions
@@
where $\mathscr{C}_{\mathfrak{k},T}^{i}:=\mathscr{C}_{\mathfrak{k},T}(A_{i})$. We stress that the stiffnes matrix contribution for a fixed element $T$ is a coarse-scale object and inexpensive to store. At the same time, we can assemble the standard FE mass matrix.
@@ 2.4.2 Online Phase
$\widehat{a}_{\mathfrak{k}}=\sum\limits_{T\in\mathcal{T}_{H}}\widehat{a}_{\mathfrak{k},T}.$
@@
The Offline-Online MLOD problem (OLOD problem): Find $(\hat{\lambda}_{H,\mathfrak{k}},\hat{u}_{H,\mathfrak{k}})\in\mathbb{R}\times V_{H}$ such that,
@@
$\hat{u}^{\mathrm{ms}}_{H,\mathfrak{k}}=\hat{u}_{H,\mathfrak{k}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\hat{u}_{H,\mathfrak{k}}\;\;\;\;\;\text{with }\;\;\;\;\widehat{\mathscr{C}}_{\mathfrak{k}}:=\sum\limits_{T\in\mathcal{T}_{H}}\sum\limits_{i=0}^{\mathcal{N}}\mu_{i}\mathscr{C}_{\mathfrak{k},T}^{i}.$
@@
Remark 2.5.
Under the considered hypercube domain structure with periodic boundary conditions, the local LOD stiffness matrix for $A_{\varepsilon}$ and the selection of offline coefficients are exactly the same for every element making the computation far more efficient and reducing the required storage. For more complex domains or different boundary conditions, the method can be adapted by modifying the offline phase accordingly. This adaptation would involve computing and storing the matrix contributions for all possible patch configurations along with their corresponding offline coefficients. If the number of such configurations remains relatively small, as in the case of a highly structured mesh and domain, the additional computational cost may still be practical.
Remark 2.6.
The possibility of using the same patch by transferring it through the whole domain contribute to a higher computational efficiency. While larger patches improve accuracy by capturing non-local interactions, they also increase storage and computational costs as the number of potential defect configurations in the offline phase grows exponentially with patch size. A recent development is the Super Localized Orthogonal Decomposition (SLOD) method proposed in [HP23], which enables the use of smaller patches without significantly compromising accuracy. While combining the offline-online strategy with the SLOD method could be attractive, our current approach is particularly based on the Petrov-Galerkin LOD formulation, and the matrix assembly is tailored to the standard LOD method. Therefore, it is not straightforward to apply our strategy to the SLOD framework in its current form.
@@
$E^{2}_{T}:=\max\limits_{v\in V_{H}:v|_{T}}\frac{\norm{\left(A^{1/2}-A^{-1/2}\overline{A}\right)\chi_{T}\nabla v-\sum\limits_{i=0}^{\mathcal{N}}\mu_{i}\left(A^{1/2}-A^{-1/2}A_{i}\right)\nabla\left(\mathscr{C}_{\mathfrak{k},T}^{i}v\right)}^{2}_{L^{2}(U_{\mathfrak{k}}(T))}}{\norm{v}^{2}_{A,T}}$
with $\overline{A}=\sum\limits_{i=0}^{\mathcal{N}}\mu_{i}A_{i}$. Then, for any $v_{1},v_{2}\in V_{H}$ it holds that
|$\absolutevalue{\left(\widehat{a}_{\mathfrak{k}}-\widetilde{a}_{\mathfrak{k}}\right)\left(v_{1},v_{2}\right)}\lesssim\mathfrak{k}^{d/2}\left(\max\limits_{T\in\mathcal{T}_{H}}E_{T}\right)\norm{v_{1}}_{\mathcal{A}}\norm{v_{2}}_{\mathcal{A}}.$ (11)
@@
$\begin{split}\norm{\left({\mathscr{C}_{\mathfrak{k}}}-\widehat{\mathscr{C}}_{\mathfrak{k}}\right){v}}^{2}_{\mathcal{A}}\lesssim\;\mathfrak{k}^{d}\left(\max_{T\in\mathcal{T}_{H}}E_{T}\right)^{2}\norm{v}^{2}_{\mathcal{A}}.\end{split}$ (12)
@@
$\begin{split}\widehat{a}_{\mathfrak{k}}(v,v)\geq\left(c_{1}-c_{2}g(\mathfrak{k})-\eta_{\mathfrak{k}}\right)\norm{v}^{2}_{\mathcal{A}}\end{split}$ (13)
@@ 2.5 An alternate offline-online strategy
In this section, we define an enhanced offline-online strategy for random checkerboard-type coefficients, particularly when $Q=[0,1]^{d}$. Note that, if $Q$ is a proper subset of $[0,1]^{d}$, the sum constraint $\sum_{i}\mu_{i}=1$ is necessary to guarantee the correct value of the coefficient $A$ at points $x$ where defects can never occur. One example is the case of random erasure and points $x$ corresponding to the value $A_{\mathrm{per}}$ independent of the realization. That means, in random erasure, the sum constraint $\sum_{i}\mu_{i}=1$ is crucial to obtain the correct multiscale basis functions that preserve the mean value of the solution over each coarse element and one cannot relax this restriction for the underlying random erasure model. However, in the random checkerboard case, the sum constraint $\sum_{i}\mu_{i}=1$ is merely used to guarantee uniqueness of the $\mu_{i}$. The alternate method we propose takes advantage of this by relaxing the sum constraint to be a real number $s$ that is adapted to the underlying distribution to obtain more accurate results. In the explained random checkerboard example, we use the Bernoulli distribution. However, other probability distributions on the interval $[\alpha,\beta]$, such as the uniform distribution, are possible in general as well. To this end, we consider the coefficient $A$ of the following form
$A(x,\omega)=\sum\limits_{j\in J}\chi_{\varepsilon(j+Q)}(x)\tilde{\beta}_{j}(\omega)$
@@
$A^{i}_{\text{harm}}=\frac{\mathcal{N}}{\left(\mathcal{N}-1\right)\left(1/\alpha\right)+(1)\left(1/\beta\right)}.$ (15)
@@
for $A_{\mathrm{per}}=\alpha$ and $B_{\mathrm{per}}=\beta-\alpha$ in the random checkerboard set-up, in which the underlying linear system of equations assigns $\tilde{\beta}_{i}$ value for each $\varepsilon$ block on the patch via the characteristic function $\chi_{\varepsilon\left(i+Q\right)}(x)$. Further observe that $A_{0}$ is nothing but the background coefficient $A_{\text{per}}$. Hence, we deduce
$\begin{split}\mu_{i}=\frac{\tilde{\beta}_{i}-\alpha s}{\beta-\alpha},\;\;i\geq 1,\hskip 56.9055pt\mu_{0}=\frac{s\left(\beta-\alpha\right)-\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}{\beta-\alpha}.\end{split}$ (16)
@@
$\begin{split}A^{\mu}_{\text{harm}}|_{T}&=\frac{\absolutevalue{T}}{\varepsilon}\frac{1}{\beta-\alpha}\Biggl{(}\frac{s(\beta-\alpha)-\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}{\mathcal{N}\left(1/\alpha\right)}+\frac{\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}{\left(\mathcal{N}-1\right)\left(1/\alpha\right)+\left(1/\beta\right)}\Biggr{)}\\ &=\frac{\absolutevalue{T}}{\beta-\alpha}\Bigg{(}\frac{s\left(\beta-\alpha\right)-\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}{\mathcal{N}\varepsilon/\alpha}+\frac{\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}{\left(\mathcal{N}-1\right)\varepsilon/\alpha+{\varepsilon}/{\beta}}\Bigg{)}\end{split}$
With this, we obtain for the error that
$\begin{split}A_{\text{harm}}|_{T}-A^{\mu}_{\text{harm}}|_{T}&=\frac{\absolutevalue{T}}{\varepsilon}\left[\frac{1}{\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}\big{)}^{-1}}-\frac{1}{\beta-\alpha}\Bigg{(}\frac{s\left(\beta-\alpha\right)-\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}{\mathcal{N}}\alpha+\frac{\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}{\frac{\left(\mathcal{N}-1\right)\beta+\alpha}{\alpha\beta}}\Bigg{)}\right]\\ &=H\big{(}\tilde{\beta}\big{)}-s\alpha+\frac{\alpha}{\beta-\alpha}\Bigg{(}\frac{\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}{\mathcal{N}}\Bigg{)}\mathcal{N}-\frac{\mathcal{N}^{2}\alpha\beta}{\left(\beta-\alpha\right)}\Biggl{[}\Biggl{(}\frac{\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}{\mathcal{N}}\Biggr{)}\frac{1}{\left(\mathcal{N}-1\right)\beta+\alpha}\Biggr{]}\\ &=H\big{(}\tilde{\beta}\big{)}-s\alpha+\frac{M\big{(}\tilde{\beta}-\alpha s\big{)}}{\left(\beta-\alpha\right)}\frac{\alpha\mathcal{N}\left(\alpha-\beta\right)}{\left(\mathcal{N}-1\right)\beta+\alpha}\\ &=H\big{(}\tilde{\beta}\big{)}-s\alpha-\alpha M\big{(}\tilde{\beta}-\alpha s\big{)}\frac{1}{\left(1-1/\mathcal{N}\right)\beta+\alpha/\mathcal{N}}\end{split}$
@@
with $H\big{(}\tilde{\beta}\big{)}=\mathcal{N}\left(\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}\big{)}^{-1}\right)^{-1}$ and $M\big{(}\tilde{\beta}-\alpha s\big{)}=\left({\sum\limits_{i=1}^{\mathcal{N}}\big{(}\tilde{\beta}_{i}-\alpha s\big{)}}\right)/{\mathcal{N}}$. Our goal is to make this error as small as possible because this will imply a small consistency error. We consider the limit $\mathcal{N}\to\infty$ which corresponds to vanishing periodicity length $\varepsilon\to 0$, and try to find $s$ such that the error vanishes in this limit. In other words, we find $s$ in such a way that in the homogenization limit, there is no consistency error and our offline-online strategy retrieves the homogenized coefficient. Denote
$\overline{H}=\lim\limits_{\mathcal{N}\to\infty}H\big{(}{\tilde{\beta}}\big{)}\;\;\;\text{ and }\;\;\;\overline{M}=\lim\limits_{\mathcal{N}\to\infty}M\big{(}{\tilde{\beta}}\big{)}.$
@@
Note that $\overline{H}$ and $\overline{M}$ can be deduced from the probability law of $\tilde{\beta}$ as certain expectations by the law of large numbers. Therefore, as $\mathcal{N}\to\infty$, we obtain
@@
$A_{\text{harm}}|_{T}-A^{\mu}_{\text{harm}}|_{T}=\overline{H}-s\alpha-\frac{\alpha}{\beta}\left(\overline{M}-\alpha s\right)=0.$
@@
$s=\frac{\beta\overline{H}-\alpha\overline{M}}{\alpha(\beta-\alpha)}.$
@@
$\overline{M}=\alpha\left(1-p\right)+\beta p\;\;\;\text{ and }\;\;\;\overline{H}=\frac{\alpha\beta}{\beta+\left(\alpha-\beta\right)p}.$
@@
$s=1+\frac{\left(\beta-\alpha\right)p^{2}}{\beta+\left(\alpha-\beta\right)p}.$
Since the alternate $s$-value deviates more from $1$ as $p$ increases, we expect the influence of our new strategy to be greater for larger $p$. This will be studied in more detail numerically in Section 5, specifically numerical experiments for values of $s$ in accordance with the Example 2.8 are illustrated in Figures 5 and 6 in Section 5 below.